------------------------------------------------------------ Date: Tue Aug 29 16:33:11 2000, Shift: 08:00-16:00 Supervisor: EJK, with: FR, JJG Log book#5 pages 41 - 43 Shift Summary: ------------------------------------------------------------ 1. RHIC changed offending component in Yellow ring overnight. Vacuum has improved by factor of 5 (says Dana). 14.00 Still high backgound seen in MTPC1. ZDC-Y rate w/clock: 3200, ZDC-B rate w/clock: 800 Coincidences: 60 cts/sek (scaler 1). The latter number is much too small for the intensities Y=25 10**9, B: 15 10**9. Would expect 300 based on counting house formula. Ask MCR to steer beam. This improves scaler 1 (coinc) to 200 cts/sek. Now MTPC1 looks much better. Decide to start run. 2. Run 2470 started at 16.00. Event rate is 100 cts/sec. Trigger 1 scale down is 1, trigger 6 scale down is 1. 3. All detectors operating OK. Polarities of D1,D2,D5 have been checked overnight with probe. They should all be the same. Nominal polarity is B. Do we also know that B IS B? MRS at 90 deg, . FS at 5 deg. ------------------------------------------------------------
This archive was generated by hypermail 2b29 : Tue Aug 29 2000 - 16:36:48 EDT